OpenStudy (anonymous):

Hello. I'm trying to get someone to provide a break down of how you get 1 1/6 as the answer for subtracting 2/3 yards of fabric from 3 1/2 yards.

OpenStudy (anonymous):

3(1/2) yards wil be 7/2 yards now 7/2 - 2/3 take lcm and subtract the numerators

OpenStudy (anonymous):

How do you get 1 1/6 from that?

OpenStudy (anonymous):

you dont get that answer

OpenStudy (anonymous):

well you should because when you do 7/2 - 2/3, you should get some number close to 2

OpenStudy (anonymous):

7/2 - 2/3 lcm of 2,3 = 6 7(3) - 2(2)/ 6 = 21-4/6 17/6 write 17/6 in mixed form